OpenStudy (anonymous):

4.A furniture maker is making a chair and needs 12 feet of wood. He has chosen to make the chair out of maple, which costs $25 per foot, and pine, which costs $28 per foot. If the furniture maker wants to spend a total of $25.75 per foot, how many feet of each type of wood will he need?

OpenStudy (anonymous):

answer plase

OpenStudy (anonymous):

somebody help

OpenStudy (anonymous):

A furniture maker is making a chair and needs 12 feet of wood. He has chosen to make the chair out of maple, which costs $25 per foot, and pine, which costs $28 per foot. If the furniture maker wants to spend a total of $25.75 per foot, how many feet of each type of wood will he need? | -------------------------- Let X = Number of FEET at $25 Let Y = Number of FEET at $25.75 X + Y = 12 25X + 25.75Y = 28 X = 12 - Y 25(12 - Y) + 25.75Y = 28 300 - 25Y + 25.75Y = 28 0.75Y = -272 Y = -362.667 X = 374.667 I HOPE I HELPED HON!